OpenStudy (anonymous):

A car is traveling at a constant velocity. After two hours the car has traveled 90 miles. After five hours the car has traveled 225 miles. Assuming a linear equation fits this situation, how far will the car have traveled after seven hours?

OpenStudy (ddcamp):

Let p(t) be the position (distance) the car has traveled after t hours. We know that p(t) is linear (i.e. p(t) = m×t) The first step is to find m. p(2) = 90 → 90 = m×2 p(5) = 225 → 225 = m×5

OpenStudy (ddcamp):

Once you have m, your new equation will be p(t) = m×t. Replace t with 7 and you should be able to find your answer.

OpenStudy (anonymous):

thanks